of ft long and 48.0 ft wide and has 8.0-ft-high cellings. What is the volume of the interior the house in cubic meters and cubic centimeters? m3 cm
Answer #1

1 ft= 0.3048 m

Therefore, 1 ft- (0.3048) 0.0283 m 7n

Hence volume of the interior of the house is,

V = 59 × 48 × 8 ft = 59 × 48 × 8 × 0.0283 m3 = 641.1648 m

y 641 . 1 648 m3-641 . 1648x 106 cm3 ~ 6.4 1 108cm